OpenStudy (anonymous):

What is the value of x? x/4-5=6

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Is this it? \(\dfrac{x}{4} - 5 = 6\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i don't get it plz help and explain it for me

OpenStudy (igreen):

Okay, first we add 5 to both sides.. \(\dfrac{x}{4} = 6 + 5\) Add that, what do you get?

OpenStudy (anonymous):

11. so x=11?

OpenStudy (anonymous):

@iGreen so x=11?

OpenStudy (igreen):

No, that gives us: \(\dfrac{x}{4} = 11\) Now we have to multiply 4 to both sides: \(x = 11 \times 4\)

OpenStudy (anonymous):

x=44?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thanks @iGreen

OpenStudy (anonymous):

x=44 because 6+5=x/4 so x=11*4